# ORBIT: An Object Property Reasoning Benchmark for Visual Inference Tasks

### v1

@@ A Framework for Object Property Evaluation
Our dataset curation follows a three-phase pipeline (Figure 2). We collect images from public sources corresponding to the three image types. Then, we ask an annotator to formulate candidate question-answer pairs for all images, inspired by AI-generated questions. Then, the questions and answers undergo additional rounds of human refinement for quality assurance. The outcome of these steps is assembled in the last phase as ORBIT, where we present statistics about the benchmark content. We detail each step below.
@@
Image Collection. To ensure diagnostic quality and mitigate data contamination risks from existing benchmarks (Dodge et al. 2021), a collection of previously unsourced images is created. We pick 15 themes grouped into three categories from the McGyver dataset (Tian et al. 2025), and add additional 11 for a total of 26 themes (see Table 7 for the complete list). The three broad categories, defined in (Tian et al. 2025), are: indoors/household (e.g., kitchen), neutral (e.g., garage), and outdoors (e.g., zoo) . We use these themes as queries for all three image types, which facilitates direct comparisons between images of comparable scenes. We collect public images by querying a variety of sources: Google Images, Unsplash, and Freepik for photographic and animated images; GPT-4o and Grok3 for AI-generated ones (example generative AI prompts are provided in Table 8). We explicitly exclude diagrammatic images (Lu et al. 2022). In total, we collect 360 images (120 per type).
@@
Formulation of Candidate QA-pairs. We formulate initial question-answer pairs using a semi-automatic procedure. Namely, we prompt an MLLM to generate three questions (one per reasoning level) following the framework defined in the previous section. The MLLM is selected at random from three closed-source models: GPT-4o (OpenAI et al. 2024), Claude 3.7 Sonnet (Anthropic 2025), and Gemini 2.0 Flash (Google et al. 2025a). The resulting 1,080 question-answer pairs are then leveraged by human annotators, instructed to revise or replace the questions and answers to ensure high quality. Namely, humans ensure that the question is precise and it fits the restrictions of our evaluation framework, as well as that the answer count is not too high (we cap the answers to a maximum of 10). A pool of five annotators annotates the images, each responsible for 72 images. When possible, the annotators are instructed to abstract away from the specific object types, e.g., to use “how many objects are made of rubber” instead of “how many tires are made of rubber?”. In practice, the annotators rewrote a vast majority of the questions, and nearly all answers, as the MLLM-generated questions were often ambiguous or repetitive, and many answers were incorrect.
Figure 3: ORBIT question composition in percentages.
@@
Quality Assurance and Consolidation. Human annotators may introduce controversial assumptions and biases. To ensure a broader consensus and catch ambiguous questions, we perform two rounds of quality assurance. First, each of the 360 images is randomly assigned to an annotator (from the same pool) who did not annotate the question-answer in the previous phase. This annotator is provided with the question and the image, and asked to provide an answer. After this round, we compared the answers by the two annotators for each question. We observed that the annotators agreed on 756 question answers (70%). We identified three key disagreement sources: 1) ambiguity of whether an object should be counted as a group or split into individual components (e.g., window versus window pane as a glass object); 2) objects that partially satisfy a property (e.g., knife that is partially made of metal or black); and 3) objects with unclear definitions (e.g., container). We resolved these disagreements by another round of quality assurance, where a third annotator (different from the first two) looked at the question, answer, and the comments, and attempted to resolve the disagreement by rephrasing the questions to make them more precise, or if that was challenging, to replace them with a new question-answer pair from the same reasoning complexity level as the original question.
@@
Benchmark Composition and Analysis. ORBIT integrates the images, questions, and final ground truth answers in a dataset containing 1,080 questions about 360 high-quality images of photographic, animated, and AI-generated types in equal proportions. Figure 3 visualizes the distribution of questions in ORBIT on the property of the object and the dimensions of the complexity of reasoning. The overall distribution across object properties is relatively balanced. We observe a larger portion of taxonomic over physical questions in the direct recognition category, and similarly, slightly more relational than functional questions in the inference category. Curiously, when creating counterfactual questions, annotators prioritized physical and taxonomic over relational and functional knowledge. We also estimated the benchmark difficulty for humans by asking two new annotators to answer 90 questions about 30 random images from ORBIT, split across reasoning levels and object properties. Their overall accuracy of 74% and an annotator agreement of 68% indicate that while the task is challenging for humans, they answer most questions correctly.
Experimental Setup
@@
Model Selection. To ensure representative evaluation, we include a variety of non-instruction-tuned and instruction-tuned open-source VLMs, as well as a fine-tuned VLM and a closed-source model. As non-instruction-tuned open-source VLMs, we include BLIP-2 (Li et al. 2023) combined with OPT-2.7b and OPT-6.7b (Zhang et al. 2022), and Fuyu-8b (Bavishi et al. 2023). We include the instruction-tuned open-source models BLIP-2 paired with Flan T5-xxl (Chung et al. 2024), Qwen2.5-VL (3B, 7B, and 32B Instruct models) (Bai et al. 2025), and InternVL3 family (8B and 14B variants) (Zhu et al. 2025) prominent on the Hugging Face OpenVLM Leaderboard (Duan et al. 2024), and Gemma 3 (Google et al. 2025b) with 27B parameters. Our study incorporates the SpatialVLM fine-tuned version of Qwen2.5-VL for spatial relations and depth estimation in VQA (Chen et al. 2024), SpaceThinker-Qwen2.5-VL, claimed by the authors to be their most accurate fine-tuned specialized reasoning model. As a closed-source VLM, we include only GPT-4o-mini due to budget considerations (OpenAI et al. 2024). Finally, we include a baseline that samples a valid answer (0-10) uniform randomly, i.e., at 9.09% accuracy. All models are evaluated in a zero-shot setting using standard hyperparameters.
@@
Answer Extraction. To extract numeric answers from model outputs, we employ regular expressions for detection and perform exact string matching against ground-truth labels. Prompts were designed to facilitate this extraction by requesting only the total count (numeric) or a count followed by object lists (see Table 6). The regular expression extraction is kept flexible to obtain counts provided in textual or Roman numeral format (observed with some non-instruction-tuned models). Counts greater than ten are uninformative in the analysis of deviation from the ground-truth and therefore are considered invalid. We interpret missing outputs as a prediction of zero.
@@
Evaluation Metrics. Following established count-based VQA benchmarks (Trott, Xiong, and Socher 2018; Acharya, Kafle, and Kanan 2018), we report multiple evaluation metrics. Our primary measure is micro accuracy (Acc). We report macro accuracy, where we average the accuracy scores for different counts, ranging from 0 to 10. We also report Root-Mean-Squared Error (RMSE) to quantify the typical deviation between predicted and ground-truth counts. To diagnose model bias, we include mean error, which reveals bias towards over- or under-counting. Furthermore, off-by-N accuracy (cumulative accuracy by tolerance) indicates the proximity of predictions to the correct count.
Results
@@
Table 2: Zero-shot results on ORBIT. The highest and second-highest model results are highlighted in bold and underlined, respectively. ($\uparrow$)/($\downarrow$) indicates that higher/lower values are better.
@@
How well can VLMs accurately count and reason about object properties? The results in Table 2 show that, for both the open- and closed-source categories, most models outperform the random baseline, though performance remains well below human accuracy (28% versus 74% on average). The high RMSE across many of the models indicates substantial variance in predictions, suggesting that precise numerical grounding is challenging, even for strong VLMs. The negative mean error values suggest consistent underestimation and a general bias towards undercounting. We noted a severely imbalanced prediction distribution in the outputs of some VLMs (see Table 2 and Figure 7). When allowing one counting error, the accuracy increases to 58% on average (to 73% for the best model), which indicates that many model errors are near-misses (see Figure 8 for in-depth analysis of the top-3 best performing models). For all models, the micro accuracy is higher than the macro accuracy. We look into this difference by analyzing the accuracy across ground truth counts in Figure 4, which shows that all models have a bias towards low, more frequent counts. In contrast, the accuracy for humans is uniform in answer counts, with no apparent frequency biases.
How does performance vary across VLMs? Accuracies vary widely across all models, between 12.87% and 39.91%. The open-source Qwen2.5-VL-Instruct (32B) is the best-performing model, with the highest exact and off-by-1 accuracy, the lowest RMSE, and a low mean error. A small gap separates it from InternVL3 (14B), with the second-highest overall accuracy, and Gemma 3 (27B), which achieves the lowest mean error. In the closed-source category, the GPT-4o-mini exhibits moderate performance, with a high RMSE and low mean error. Table 3 presents results across reasoning levels, property dimensions, and image types, where the open-source Qwen2.5-VL-Instruct (7B) and (32B) variants demonstrate the strongest performance across nearly all categories. Overall, we observe that GPT-4o-mini performs consistently worse than most instruction-tuned open-source models. We also note a typical increase in performance with increasing model size within model families, though this may not always hold across architectures.
@@
Model Reasoning Complexity Object Properties Image Types Direct Recognition Property Inference Counter- factual Physical Taxonomic Functional Relational Photographic Animated AI Gen. BLIP-2 OPT (2.7B) 10.83 13.06 14.72 11.48 14.41 10.91 13.78 9.17 19.44 10.00 BLIP-2 OPT (6.7B) 16.67 18.06 15.28 16.39 16.52 14.09 19.08 14.17 21.11 14.72 Fuyu (8B) 18.33 22.78 18.61 17.21 17.42 19.55 25.44 16.94 23.89 18.89 BLIP-2 Flan (11B) 18.33 18.61 10.83 14.34 16.82 13.64 18.02 11.67 20.00 16.11 Qwen2.5 (3B) 26.11 29.17 19.44 21.31 29.43 17.73 28.27 20.28 27.78 26.67 Qwen2.5 (7B) 42.78 44.17 29.17 35.66 42.34 32.73 41.70 32.22 40.83 43.06 Qwen2.5 (32B) 39.17 43.06 37.50 33.61 44.44 38.18 41.34 31.94 45.00 42.78 SpaceThinker (3B) 34.44 30.28 24.17 25.00 37.24 19.55 32.51 25.56 31.11 32.22 InternVL3 (8B) 38.89 41.94 31.94 33.20 42.34 26.82 44.17 28.89 43.61 40.28 InternVL3 (14B) 42.78 42.22 34.17 30.33 48.05 36.82 40.28 32.22 48.61 38.33 Gemma3 (27B) 33.89 35.56 25.83 26.23 34.53 28.64 35.69 27.78 36.39 31.11 GPT-4o mini 28.89 33.61 28.61 21.31 34.23 25.00 37.81 24.44 38.89 27.78 Model average 29.34 31.04 24.18 23.83 31.48 23.63 31.50 22.94 33.05 28.49 Human 1 73.33 80.00 73.33 58.33 92.00 71.43 80.00 63.33 83.33 80.00 Human 2 66.67 76.67 73.33 58.33 80.00 76.19 75.00 60.00 83.33 73.33 Human average 70.00 78.33 73.33 58.33 86.00 73.81 77.50 61.66 83.33 76.66
@@
Table 3: Accuracy results for each model on ORBIT across reasoning levels, object property dimensions, and image types. The highest and second highest model results are highlighted in bold and underlined, respectively.
@@
How does performance vary across reasoning levels? Based on the presented framework, we expect the progression of reasoning complexity as recognition << property inference << counterfactual. However, the results among the top-tier VLMs (Qwen2.5-VL-Instruct, InternVL3, Gemma 3, and GPT-4o-mini) suggest a progression property inference << recognition << counterfactual. This suggests that modern VLMs still exhibit poor low-level perceptual capabilities, which aligns with several recent findings. First, Jiang et al. (2024) showed poor perceptual performance of MLLMs for fine-grained visual analysis tasks, including counting. Second, Zhang et al. (2024a) found blind spots in MLLMs with sensitivity to visual quality and distractors in the image. And, finally, Yuan et al. (2021) showed perception failures in the MetaVQA framework and highlighted the need for detailed perceptual grounding of VLMs. The lower performance on counterfactual questions is indicative of the difficulty with changes and hypotheticals and their reliance on commonsense reasoning and learned associations, areas where modern VLMs lack better grounding (Li et al. 2024). Similarly to VLMs, humans perform best on inference questions, but struggle less with counterfactual reasoning.
@@
How does performance vary across object properties? Among the object property dimensions, most models (as well as humans) have a stronger grasp on taxonomic and relational questions compared to physical and functional ones. This finding is in line with recent work that points out the struggle of LLMs with physical object attributes (Wang et al. 2023), physical interactions (Aroca-Ouellette et al. 2021), functional requirements (Qasemi et al. 2022), and atypical affordances (Tian et al. 2025). We note an improvement obtained by the fine-tuned SpaceThinker-Qwen2.5-VL (3B) model compared to its non-fine-tuned counterpart, Qwen2.5-VL (3B), across the four object property dimensions. This shows that besides relational questions, the other object properties also benefit from spatial tuning of Qwen2.5 (3B). We also observe that the InternVL3 models have a better grasp of taxonomic and relational questions compared to other models. The Qwen2.5-VL (7B) and (32B) model variants demonstrate consistent performance across all property types, with particular strength in physical and functional questions compared to others. In contrast, the non-instruction-tuned models face difficulty across all properties, stressing their weakness in visual object reasoning.
@@
How does performance vary across image types? We observe a consistently higher performance for AI-generated and animated images, compared to photographic images. We hypothesize that this is due to the visually cleaner and less cluttered nature of AI-generated and animated images compared to the complex and noisy real-world scenes in photographic images. This finding aligns with those of previous benchmarks (Johnson et al. 2017; Hudson and Manning 2019). Moreover, we note that animated images are easier for most models than AI-generated images, possibly because AI-generated images may have poor focus and may include implausible objects, such as a half-cut glass floating in the air. The trend between the image difficulty animated << AI-generated << photographic is also clear in humans, which indicates that questions about photographic images may be hardest and possibly contain remaining ambiguity.
Can densely annotated datasets be used to automatically scale up ORBIT? Visual Genome (VG) (Krishna et al. 2017) is a popular dataset with high-quality structured connections between image regions, language descriptions, and WordNet (Miller 1995) synsets. To investigate whether the annotations of VG can support the generation of large-scale instantiations of ORBIT, we devised a procedure that generates taxonomic questions by manipulating the WordNet synsets. This experiment exposed three limitations: lack of contextualization of classes (e.g., kitchen utensil instead of a bowl in Figure 11-left), incorrect questions or answers (e.g., about uncountable objects like grass in Figure 11-right), and classes in the WordNet hierarchy that are not overly specific (e.g., derby horse race) or generic (e.g., entity). Thus, while every image in VG enables many taxonomic questions to be created, it remains challenging to automatically distinguish high- from low-quality question-answer pairs. More fundamentally, creating question-answer pairs for other object properties (e.g., functional) and reasoning types (e.g., counterfactual) is hindered by the absence of annotation. While LLMs can be employed to generate questions for these categories, we established that these questions and answers are often incorrect or ambiguous (e.g., suggesting that buildings, doors, and windows can all provide shelter). In conclusion, we find that densely annotated image datasets remain insufficient to generate high-quality large-scale versions of ORBIT. Yet, they could potentially be used to speed up human annotation by providing an initial set of samples or for distant supervision of VQA models.
Conclusions

### v2

@@ 4 Benchmark Construction
Our dataset curation follows a four-phase pipeline (Figure 2). We collect images from public sources corresponding to the three image types. Then, we ask an annotator to formulate candidate question-answer pairs for all images, inspired by AI-generated questions. Then, the questions and answers undergo additional rounds of human refinement for quality assurance, which results in OPTICS-CNT. OPTICS-CMP is then systematically constructed from samples in OPTICS-CNT. Finally, we present statistics about the two benchmarks. We detail each step below.
@@
Image Collection. To ensure diagnostic quality and mitigate data contamination risks from existing benchmarks (Dodge et al., 2021), we collect previously unsourced images. We define 26 domains grouped into three categories based on the MacGyver dataset (Tian et al., 2025) (see Table 6 for details). We use these domains as queries across all three image types, enabling fair comparisons of images from comparable scenes. We collect public images by querying a variety of sources: Google Images, Unsplash, and Freepik for photographic and animated images; GPT-4o and Grok3 for AI-generated ones (see Table 5 for example generative AI prompts). We explicitly exclude diagrammatic images (Lu et al., 2022). In total, we collect 360 images (120 per type).
@@
Formulation of Candidate QA-pairs. We formulate initial question-answer pairs using a semi-automatic procedure. Namely, we prompt a multimodal large-language model (MLLM) to generate three counting questions (one per reasoning level) following the framework defined in the previous section. The MLLM is selected at random from GPT-4o (OpenAI et al., 2024), Claude 3.7 Sonnet (Anthropic, 2025), and Gemini 2.0 Flash (Google et al., 2025a). The resulting 1,080 question-answer pairs are then used by five human annotators, who are instructed to revise or replace them to ensure high quality. Namely, humans ensure that the question is precise and fits the restrictions of our evaluation framework, and that the answer count is not too high (we cap the answers at 10). Each of the annotators is responsible for 72 images. When possible, the annotators are instructed to abstract over the specific object types, e.g., prioritizing “how many objects are made of rubber?” over “how many tires are made of rubber?”. In practice, the annotators rewrote a vast majority of the questions and nearly all answers, as many MLLM-generated questions were ambiguous or repetitive, and most answers were incorrect.
@@
Quality Assurance and Consolidation of OPTICS-CNT. Human annotators may introduce controversial assumptions and biases. To ensure a broader consensus and catch ambiguous questions, we perform two rounds of quality assurance. First, each of the 360 images is randomly assigned to an annotator (from the same pool) who did not annotate the QA pair in the previous phase. This annotator is provided with the question and the image, and asked to provide an answer. Comparing the two annotators’ answers for each question, we observed that they agreed on 756 answers (70%). We identified two key disagreement sources: 1) Semantic ambiguity: Uncertainty in the definition, including objects that partially satisfy a property (e.g., a knife that is partly metal) or objects with unclear definitions (e.g., a container). 2) Visual ambiguity: Uncertainty in the visual information, such as deciding whether to count a partially visible chair. We addressed these disagreements by another round of quality assurance, where a third, different annotator looked at the question, answer, and comments and attempted to resolve the disagreement by rephrasing the questions to make them more precise, or if that was challenging, to replace them with a new QA pair from the same reasoning complexity level. This results in OPTICS-CNT.
Construction of OPTICS-CMP. To introduce an evaluation axis beyond exact counting and diversify question types, we devise OPTICS-CMP, which assesses the ability to make relative judgments between images. OPTICS-CMP is derived from annotated VQA samples in OPTICS-CNT. We first group questions by object properties and reasoning complexity, then pair them using semantic similarity Reimers and Gurevych (2019). For each pair, the corresponding images are resized and concatenated. Then, a comparison question is generated using the template in Figure˜3, and an answer is automatically derived from the OPTICS-CNT labels (see example in Figure˜9).
Let AA be the answer to left question for the left image. Let $B$ be the answer to right question for the right image. Which is greater, AA or $B$?
Figure 3: OPTICS-CMP question template.
Figure 4: OPTICS-CNT question composition.
@@
Benchmark Composition and Analysis. OPTICS-CNT integrates the images, questions, and final ground truth answers in a dataset containing 1,080 questions about 360 high-quality images of photographic, animated, and AI-generated types in equal proportions. The overall distribution across object properties is also relatively balanced (Figure 4). We observe a larger portion of taxonomic over physical questions in the direct recognition category, and similarly, slightly more relational than functional questions in the inference category. Curiously, when creating counterfactual questions, annotators prioritized physical and taxonomic over relational and functional knowledge. Meanwhile, OPTICS-CMP comprises 2,122 comparison questions created from merged image pairs, where the count difference spans from 1 to 9. Notably, 80% of the comparisons fall within a difference of 1–4. The most common question pairs in OPTICS-CMP involve direct recognition of an object’s taxonomic (56.7%) or physical (22.1%) properties (details in Figure˜10). We also estimated the OPTICS-CNT difficulty for humans by asking two new participants to answer 90 questions about 30 random images from OPTICS-CNT, split across reasoning levels and object properties. Their overall accuracy of 74% indicates that while the task is challenging for humans, they answer most questions correctly.
5 Experimental Setup
@@
Model Selection. To ensure representative evaluation, we include a variety of non-instruction-tuned and instruction-tuned open-source VLMs, as well as a fine-tuned VLM and a closed-source model. As non-instruction-tuned open-source VLMs, we include BLIP-2 (Li et al., 2023) combined with OPT-2.7b and OPT-6.7b (Zhang et al., 2022), and Fuyu-8b (Bavishi et al., 2023). We include the instruction-tuned open-source models BLIP-2 paired with Flan T5-xxl (Chung et al., 2024), Qwen2.5-VL (3B, 7B, and 32B Instruct models) (Bai et al., 2025), and InternVL3 family (8B and 14B variants) (Zhu et al., 2025) prominent on the Hugging Face OpenVLM Leaderboard (Duan et al., 2024), and Gemma 3 (Google et al., 2025b) with 27B parameters. Our study incorporates the SpatialVLM fine-tuned version of Qwen2.5-VL for spatial relations and depth estimation in VQA (Chen et al., 2024), SpaceThinker-Qwen2.5-VL, claimed by the authors to be their most accurate fine-tuned specialized reasoning model. As a closed-source VLM, we include only GPT-4o-mini due to budget considerations (OpenAI et al., 2024). All models are evaluated in a zero-shot setting using standard hyperparameters.
@@
Answer Extraction. For both benchmarks, our VLM prompts facilitate answer extraction by requesting either the total count or a count followed by objects for OPTICS-CNT (see Table˜4), and only “A” or “B” to indicate which value is greater for OPTICS-CMP. To extract answers from OPTICS-CNT (numeric label) and OPTICS-CMP (A or B label), we employ flexible regular expressions for detection and perform exact string matching against ground-truth labels. For OPTICS-CNT, we interpret missing outputs as a prediction of zero, and though during data curation we restrict object counts to 10 per image, we allow model outputs with counts greater than 10 to enable analysis of deviation from ground truth.
@@
Evaluation Metrics. Following established count-based VQA benchmarks (Trott et al., 2018; Acharya et al., 2018), we report multiple evaluation metrics for OPTICS-CNT. Our primary measure is micro accuracy (Acc). We report macro accuracy, which averages the accuracy scores across different counts, ranging from 0 to 10. We also report Root-Mean-Squared Error (RMSE) to quantify the typical deviation between predicted and ground-truth counts. To diagnose model bias, we include mean error, which reveals bias towards over- or under-counting. Furthermore, off-by-N accuracy (cumulative accuracy by tolerance) indicates the proximity of the predictions to the correct count. We also report accuracy per count difference for OPTICS-CMP, expecting an increase in accuracy for larger differences.
6 Results
@@
Table 2: Zero-shot results on OPTICS-CNT. The highest and second-highest model results are highlighted in bold and underlined, respectively. ($\uparrow$)/($\downarrow$) indicates that higher/lower values are better. ($\rightarrow 0$) signifies values closer to zero are better. means the result is attributed to inductive bias.
How well can VLMs accurately count and reason about object properties? The results for OPTICS-CNT in Table 2 show that most models perform above random, but well below human accuracy (28% versus 74% on average). The high RMSE across many of the models indicates substantial variance in predictions, suggesting that precise numerical grounding is challenging, even for strong VLMs. The negative mean error values suggest consistent underestimation and a general bias towards undercounting. We note a severely imbalanced prediction distribution in the outputs of some VLMs (detailed in Figure 14). When allowing for one counting error, the accuracy increases to 58% on average (to 73% for the best model), which indicates that many model errors are near-misses (see Figure 15 for an in-depth analysis of the top-3 best performing VLMs). For all VLMs, the micro accuracy is higher than the macro accuracy because all models exhibit a bias towards lower, more frequent counts, as shown in Figure 5. In contrast, human accuracy is uniform across answer counts, with no apparent frequency bias. We further analyze the sources of errors for both humans and VLMs at the end of this section.
How does counting performance vary across VLMs? Accuracies vary widely across models, ranging from 12.87% to 39.91%. The open-source Qwen2.5-VL-Instruct (32B) is the best-performing model, with the highest exact and off-by-1 accuracies, the lowest RMSE, and the lowest mean error. A small gap separates it from InternVL3 (14B), with the second-highest overall accuracy, and Gemma 3 (27B), which achieves the lowest mean error. The closed-source GPT-4o-mini exhibits moderate performance, with a high RMSE and low mean error. We also note a typical increase in performance with increasing model size within model families, though this may not always hold across architectures. Table 3 shows that Qwen2.5-VL-Instruct (7B) and (32B) demonstrate the strongest performance across nearly all reasoning levels, property dimensions, and image types, while GPT-4o-mini performs consistently worse than most instruction-tuned open-source models.
@@
Model Reasoning Complexity Object Properties Image Types Direct Recognition Property Inference Counter- factual P T F R Photo Animated AI Gen. BLIP-2 OPT (2.7B) 10.83 13.06 14.72 11.48 14.41 10.91 13.78 9.17 19.44 10.00 BLIP-2 OPT (6.7B) 16.67 18.06 15.28 16.39 16.52 14.09 19.08 14.17 21.11 14.72 Fuyu (8B) 18.33 22.78 18.61 17.21 17.42 19.55 25.44 16.94 23.89 18.89 BLIP-2 Flan (11B) 18.33 18.61 10.83 14.34 16.82 13.64 18.02 11.67 20.00 16.11 Qwen2.5 (3B) 26.11 29.17 19.44 21.31 29.43 17.73 28.27 20.28 27.78 26.67 Qwen2.5 (7B) 42.78 44.17 29.17 35.66 42.34 32.73 41.70 32.22 40.83 43.06 Qwen2.5 (32B) 39.17 43.06 37.50 33.61 44.44 38.18 41.34 31.94 45.00 42.78 SpaceThinker (3B) 34.44 30.28 24.17 25.00 37.24 19.55 32.51 25.56 31.11 32.22 InternVL3 (8B) 38.89 41.94 31.94 33.20 42.34 26.82 44.17 28.89 43.61 40.28 InternVL3 (14B) 42.78 42.22 34.17 30.33 48.05 36.82 40.28 32.22 48.61 38.33 Gemma3 (27B) 33.89 35.56 25.83 26.23 34.53 28.64 35.69 27.78 36.39 31.11 GPT-4o mini 28.89 33.61 28.61 21.31 34.23 25.00 37.81 24.44 38.89 27.78 Model average 29.34 31.04 24.18 23.83 31.48 23.63 31.50 22.94 33.05 28.49 Human 1 73.33 80.00 73.33 58.33 92.00 71.43 80.00 63.33 83.33 80.00 Human 2 66.67 76.67 73.33 58.33 80.00 76.19 75.00 60.00 83.33 73.33 Human average 70.00 78.33 73.33 58.33 86.00 73.81 77.50 61.66 83.33 76.66
@@
Table 3: Accuracy results for each model on OPTICS-CNT across reasoning levels, object property dimensions (Physical, Taxonomic, Functional, Relational), and image types. The highest and second highest model results are highlighted in bold and underlined, respectively.
@@
How does counting performance vary across reasoning levels? We expect the progression of reasoning complexity recognition << property inference << counterfactual3) to be reflected in corresponding trends in model performance. However, the results among the top-tier VLMs (Qwen2.5-VL-Instruct, InternVL3, Gemma 3, and GPT-4o-mini) reveal a progression property inference << recognition << counterfactual. This suggests that modern VLMs still exhibit poor low-level perceptual capabilities, which aligns with several recent findings: 1) Jiang et al. (2024) showed poor perceptual performance of MLLMs for fine-grained visual analysis tasks, including counting; 2) Zhang et al. (2024a) identified blind spots in MLLMs, including sensitivity to visual quality and image distractors; 3) Yuan et al. (2021) showed perception failures in the MetaVQA framework and highlighted the need for detailed perceptual grounding of VLMs. The lower accuracy on counterfactual questions is indicative of difficulty with changes and hypotheticals, reliance on learned associations, and lack of reliable grounding Li et al. (2024). Humans also perform best on inference questions, but struggle less than VLMs with counterfactual reasoning.
@@
How does counting accuracy vary across object properties? Among the object property dimensions, most models (as well as humans) have a stronger grasp of taxonomic and relational questions than of physical and functional ones. This finding is in line with recent work that points out the struggle of LLMs with physical object attributes (Wang et al., 2023), physical interactions (Aroca-Ouellette et al., 2021), functional requirements (Qasemi et al., 2022), and atypical affordances (Tian et al., 2025). We note an improvement for the fine-tuned SpaceThinker-Qwen2.5-VL (3B) model compared to its non-fine-tuned counterpart, Qwen2.5-VL (3B), across the four object property dimensions. This shows that, in addition to relational questions, other object properties also benefit from spatial tuning of Qwen2.5 (3B). We also observe that the InternVL3 models have a better grasp of taxonomic and relational questions compared to other models. The Qwen2.5-VL (7B) and (32B) variants demonstrate consistent performance across all properties, with particular strength in physical and functional questions. In contrast, the non-instruction-tuned models struggle across all properties, underscoring their weakness in visual object reasoning.
@@
How does counting performance vary across image types? We observe a consistently higher accuracy for AI-generated and animated images, compared to photographic images. We hypothesize that this is due to the visually cleaner, less cluttered nature of AI-generated and animated images compared to the complex, noisy real-world scenes in photographic images. This finding aligns with previous studies reporting similar gaps between synthetic and real-world data Johnson et al. (2017); Hudson and Manning (2019). Moreover, we note that animated images are easier for most models than AI-generated images, possibly because the latter may have poor focus and include implausible objects, such as a half-cut glass floating in the air. The trend between the image difficulty animated << AI-generated << photographic is also clear in humans, which indicates that questions about photographic images may be the hardest and possibly contain remaining ambiguity.
How well do VLMs answer comparison questions, and is this sensitive to the count differences? Based on results from OPTICS-CNT, we evaluate only the instruction-tuned models on OPTICS-CMP. This experiment confirms that VLMs struggle with counting objects, even in this relative setting. Namely, VLMs’ accuracy remains only slightly higher than random, at 58% (details in Table 9). Qwen2.5-VL (32B) remains the overall best-performing model, followed by InternVL3 (8B), which outperforms its larger (14B) variant. This pattern does not hold within the Qwen2.5-VL family, where we instead observe a steady increase in performance as model size increases. Our hypothesis that VLMs would show higher gains on larger count differences is examined in Figure 6. While some models clearly follow this trend, we note that the model performance across the count differences remains broadly stable. These findings confirm those from the absolute-counting task but also demonstrate that the failure modes of humans and VLMs may differ. We next dive deeper into the causes for errors in both humans and VLMs.
123456-940405050606070708080Ground Truth Count DifferenceAccuracy (%)Gemma27BGpt4ominiIntern14BIntern8BQwen32BQwen3BQwen7B
Figure 6: Accuracy per count difference (1-9) for instruction-tuned VLMs on OPTICS-CMP. Counts (6-9) are grouped as one.
Figure 7: Examples of errors made by humans and the model. Left: Both the model and the human made a perception error. Right: The human made a semantic ambiguity error, while the model made an unnatural error.
How similarly do humans and VLMs make mistakes? To investigate this, we selected 25 questions for which both the best VLM (Qwen2.5-VL-32b) and at least one human judge gave an incorrect answer relative to the ground truth. We defined four categories of errors: 1) Semantic ambiguity errors caused by unclear semantic constraints in the question, e.g., should an object that is half wood and half metal be counted as “metal”? 2) Visual ambiguity errors caused by ambiguous visual information in the image, e.g., should we count a chair if only part of its legs is visible and not the whole object? 3) Perception errors - honest mistakes due to overlooking or misperceiving information, e.g., not noticing a small object or miscounting the number of items. 4) Unnatural errors that are highly unlikely for a human to make, e.g., counting a glass as a metal object or counting an object that clearly does not exist.
We assigned each incorrect answer from the model and the human judge to one or more of these four categories. We observe the following (full results in Table 10): a) Humans do not make unnatural mistakes (0/25), whereas the model makes many of them (9/25). This is the most significant difference between humans and the model, which may stem from hallucination/illusion-driven errors in models, as evidenced by diagnostic suites (Guan et al., 2024). An example is shown in Figure 7 (right), where the human counted all objects that could be considered metal, while the model counted items like the glass and the cutting board, which are clearly not metal. b) For humans, the majority of mistakes (16/25) come from semantic ambiguity, not knowing whether something should be counted or how to interpret a specific condition. This finding is consistent with our analysis of image types, which concluded that photographic images may be most challenging due to remaining ambiguity in the questions or in the visual content. c) The second most common source of human error is perception errors (9/25), often due to not examining the image carefully enough and missing objects. Perception errors are similarly frequent in humans and the model (9 and 8, respectively). An example is shown in Figure 7 (left), where both the human and the model made the same mistake: they missed the turtle inside the water.
7 Conclusions